A 1.75 kilowatt winch is used to raise loads of bricks up a 24.0 meter building. If the bricks need to get to the top of the building in 2.00 minutes, what is the maximum allowable mass of a load of bricks?
1
Expert's answer
2020-11-20T07:13:52-0500

By the definition of the power, we have:


"P=\dfrac{W}{t}=\dfrac{mgh}{t}."

From this formula we can find the maximum allowable mass of a load of bricks:


"m=\dfrac{Pt}{gh},""m=\dfrac{1.75\cdot 10^3\ W\cdot 2\cdot 60\ s}{9.8\ \dfrac{m}{s^2}\cdot 24.0\ m}=893\ kg."

Answer:

"m=893\ kg."